# PEFT-as-an-Attack! Jailbreaking Language Models during Federated Parameter-Efficient Fine-Tuning

### v1

@@ IV-A Learning Effectiveness of FedPEFT Methods
IV-B Comparison of FedPEFTs under PaaA
@@
Figure 3: ASR comparison of FedPEFT methods under varying numbers of malicious clients (0, 1, and 5) across different PLMs over 20 communication rounds. Initially at Round 0, all methods show low jailbreak risk ($ASR<4\%$). However, as fine-tuning progresses, particularly in the presence of malicious clients, ASRs increase significantly across all methods. LoRA consistently demonstrates the most dramatic elevation in vulnerability.
@@
IV-C Defense with RASs
@@
Table II presents the utility (MedQA/MetaMathQA accuracy) and safety (AdvBench/JailbreakBench ASR) metrics across different RASs. Regarding utility, we observe that all RASs maintain or improve performance compared to the base model (initial accuracies are 56.9% and 80.0% for MedQA and MetaMathQA respectively) in their respective IID settings, even in the presence of malicious clients. Notably, when trained on MedQA-only data, the model achieves significant improvements in MedQA accuracy (up to 64.7% accuracy) but suffers a decline in MetaMathQA performance (down to 71.1%). Similarly, MetaMathQA-only training improves MetaMathQA accuracy (up to 83.6%) with significant performance degradations in MedQA. More importantly, the results illustrate the vulnerability of RASs: Traditional RASs (Median and GeoMed) show consistently poor defense performance ($ASR>30.7\%$) regardless of data distribution settings. While DnC and ClippedClustering effectively defend against jailbreak attacks under the MedQA setting ($ASR\leq 2\%$), they fail to maintain model safety when data is heterogeneously distributed across clients ($ASR>80\%$).
IV-D Defense with PPSA
Figure 4: Evaluation of jailbreak attacks, normal operations, and alignment processes in FedPEFT across communication rounds. This study demonstrates that both jailbreak attacks and alignment can be achieved within very few communications (less than 5). However, alignment improvements may come at the expense of reduced task performance. The experiments involved fine-tuning the three PLMs on MedQA over 14 communication rounds using LoRA. During Rounds 0–5, three malicious clients were activated, nine fine-tuning clients operated between Rounds 0–11, and three alignment clients were introduced in Rounds 11–14.
@@ V Related Work
In this section, we review related work under the context of FedPEFT in Section V-A, followed by fine-tuning-based jailbreak attacks and defenses during the adaptation of PLMs in Section V-B.
V-A Federated Parameter-Efficient Fine-Tuning
@@
Originating from the fine-tuning practices of transformer models [14, 15, 16], FedPEFT is a training paradigm aiming to adapt pre-trained models based on PEFT and FL techniques, simultaneously preserving data privacy and minimizing communication costs and computational overhead [61, 25]. Some studies have demonstrated their efficiency in training and evaluating the comparative advantages and disadvantages of different PEFT methods in terms of performance [62, 63], resource efficiency [64, 65], and model personalization [66, 67]. Some other studies focus on exploring the potential of FedPEFT techniques within resource heterogeneity FL environments, such as using heterogeneous LoRA ranks and sparsity to facilitate limited and heterogeneous system capabilities of clients [68, 69, 70].
@@
Additionally, recent work has been focusing on the security aspect, which involves identifying security threats to this emerging paradigm [71] and developing defense mechanisms to safeguard the fine-tuning process [72, 38], aiming to ensure security and maintain the performance advantage simultaneously.
V-B Fine-Tuning-based Jailbreak Attacks and Defenses
@@
In the past few years, a branch of research has been focusing on the susceptibility of PLMs to fine-tuning-based jailbreak attacks [27], where malicious actors exploit vulnerabilities in the model’s architecture or alignment to elicit the harmful behaviors. Specifically, several pioneering works [31, 73, 74] showed that PLMs aligned by RLHF or SFT can be jail-broken through fine-tuning on toxic data. Moreover, a line of work [75, 76, 77, 78, 79] focused on accessing and analyzing how jailbreak attacks disrupt the safety guardrails and the associated implications.
@@
To mitigate these vulnerabilities, researchers have been actively exploring defense mechanisms across different model training and adaptation stages. 1) Alignment stage defense aims to enhance the model’s immunization ability to fine-tune by modifying the training procedure in the alignment stage [42]. For instance, Vaccine [42] vaccinates the model by adding embedding perturbation in the alignment stage, and RepNoise improve the robustness by enforcing the representation of the harmful data to be a random Gaussian noise. 2) Fine-tuning stage defense focuses on maintaining the alignment knowledge while training on the downstream dataset [80, 81, 82]. For instance, LDIFS [80] introduces a regularizer to enforce the iterate’s embedding to be close to that of the aligned model. VLGuard [81] mixes the alignment data with the user data to train in the user finetuning stage. Lisa [82] alternatively optimizes the alignment data and the user fine-tuning data and uses a proximal regularizer to enforce proximity between iterates. 3) Post-fine-tuning stage defense adopts safety alignment techniques to restore safety guardrails after malicious fine-tuning [38, 32, 83].
@@
To the best of our knowledge, only one preliminary study [38] has investigated jailbreak attacks in the federated fine-tuning settings, wherein the authors proposed a PPSA based on a pipeline that automatically generates defense data and performs safety alignment on the server-side. However, [38] is mainly limited in the single focus on LoRA fine-tuning with a LLaMA-2 model. In contrast, our work extensively investigates several PEFT methods, PLMs, and defenses under different FL settings. Through the broader scope, we draw more generalizable conclusions about the attack’s effectiveness across various scenarios, demonstrating the limitations of current defenses and emphasizing the need for advanced approaches.
VI Conclusions and Future Work
@@
In this paper, we introduced PEFT-as-an-Attack (PaaA), a novel security threat to FedPEFT, demonstrating that PEFT methods can be exploited to bypass safety alignment and generate harmful content in response to malicious prompts. Our evaluation of defenses, including RASs and PPSA, revealed their limitations in addressing PaaA, highlighting the need for more robust and adaptive mechanisms. In the future, we will focus on developing advanced PPSA techniques for FedPEFT to balance model utility and safety. Another promising direction is integrating safety alignment directly during fine-tuning to dynamically mitigate emerging vulnerabilities without compromising model performance.
@@ Acknowledgment
The computationally intensive experiments presented in this research were performed on the Alviscluster at Chalmers University of Technology, funded by the Swedish National Infrastructure for Computing (SNIC).

### v2

@@ V-A Learning Effectiveness of FedPEFT Methods
V-B Comparison of FedPEFTs under PaaA
@@
Figure 4: ASR comparison of FedPEFT methods under varying numbers of malicious clients (0, 1, and 5) across different PLMs over 20 communication rounds. Initially at Round 0, all methods show a low jailbreak risk ($ASR<4\%$). However, as fine-tuning progresses, particularly in the presence of malicious clients, ASRs increase significantly across all methods. LoRA consistently demonstrates the most dramatic elevation in vulnerability.
@@
V-C Defense with RASs
@@
Table II presents the utility (MedQA/MetaMathQA accuracy) and safety (AdvBench/JailbreakBench ASR) metrics across different RASs. Regarding utility, we observe that all RASs maintain or improve performance compared to the base model (initial accuracies are 56.9% and 80.0% for MedQA and MetaMathQA respectively) in their respective IID settings, even in the presence of malicious clients. Notably, when trained on MedQA-only data, the model significantly improves MedQA accuracy (up to 64.7% accuracy) but declines in MetaMathQA performance (down to 71.1%). Similarly, MetaMathQA-only training improves MetaMathQA accuracy (up to 83.6%) with significant performance degradations in MedQA. More importantly, the results illustrate the vulnerability of RASs: Traditional RASs (Median and GeoMed) show consistently poor defense performance ($ASR>30.7\%$) regardless of data distribution settings. While DnC and ClippedClustering effectively defend against jailbreak attacks under the MedQA setting ($ASR\leq 2\%$), they fail to maintain model safety when data is heterogeneously distributed across clients ($ASR>80\%$).
V-D Defense with PPSA
Figure 5: Evaluation of jailbreak attacks, normal fine-tuning, and alignment processes in FedPEFT across communication rounds. This study demonstrates that both jailbreak attacks and alignment can be achieved within very few communications (less than 5). However, alignment improvements may come at the expense of reduced task performance. The experiments involved fine-tuning the three PLMs on MedQA over 14 communication rounds using LoRA. During Rounds 0–5, three malicious clients were activated, nine fine-tuning clients operated between Rounds 0–11, and three alignment clients were introduced in Rounds 11–14.
@@ VI Related Work
In this section, we review related work under the context of FedPEFT in Section VI-A, followed by fine-tuning-based jailbreak attacks and defenses during the adaptation of PLMs in Section VI-B.
VI-A Federated Parameter-Efficient Fine-Tuning
@@
Originating from the fine-tuning practices of transformer models [13, 14], FedPEFT is a training paradigm aiming to adapt pre-trained models based on PEFT and FL techniques, while preserving data privacy and minimizing communication costs and computational overhead [61]. Some studies have demonstrated their efficiency in training and evaluating the comparative advantages and disadvantages of different PEFT methods in terms of performance [62, 63], resource efficiency [64], and model personalization [65]. Some other studies focus on exploring the potential of FedPEFT techniques within resource heterogeneity FL environments, such as using heterogeneous LoRA ranks and sparsity to facilitate limited and heterogeneous system capabilities of clients [66, 67].
@@
Furthermore, recent work has also focused on the security aspect, which involves identifying security threats to this emerging paradigm [68] and developing defense mechanisms to safeguard the fine-tuning process [69, 40], with the objective of ensuring security and maintaining performance advantage simultaneously.
VI-B Fine-Tuning-based Jailbreak Attacks and Defenses
@@
In the past few years, a branch of research has concentrated on the susceptibility of PLMs to fine-tuning-based jailbreak attacks [24], where malicious users exploit fine-tuning techniques to circumvent the established safety guardrails of PLMs, hereby inducing harmful behaviors. Specifically, several pioneering works [28, 70, 71] showed that PLMs aligned by RLHF or SFT can be jailbroken by fine-tuning with toxic data. Moreover, a line of work [72, 73, 74, 75] focused on accessing and analyzing how jailbreak attacks disrupt the safety guardrails and the associated implications.
@@
To mitigate these vulnerabilities, researchers have been actively exploring defense mechanisms across different model training and adaptation stages. 1) Alignment stage defense aims to enhance the model’s immunization ability to fine-tune by modifying the training procedure in the alignment stage [44]. For instance, Vaccine [44] vaccinates the model by adding embedding perturbation in the alignment stage, and RepNoise improves the robustness by enforcing the representation of the harmful data to be a random Gaussian noise. 2) Fine-tuning stage defense focuses on maintaining the alignment knowledge while training on the downstream dataset [76, 77, 78]. For example, LDIFS [76] introduces a regularizer to enforce the embedding of the iterate to be close to that of the aligned model. VLGuard [77] mixes the alignment data with the user data to train in the user fine-tuning stage. Lisa [78] alternatively optimizes the alignment data and the user fine-tuning data and uses a proximal regularizer to enforce proximity between iterates. 3) Post-fine-tuning stage defense adopts safety alignment techniques to restore safety guardrails after malicious fine-tuning [40, 29, 79].
@@
To our knowledge, only one preliminary study [40] has investigated jailbreak attacks in federated fine-tuning settings, in which the authors proposed a PPSA based on a pipeline that automatically generates defense data and performs safety alignment on the server side. However, [40] is mainly limited in the single focus on LoRA fine-tuning with a LLaMA-2 model. In contrast, our work extensively investigates several PEFT methods, PLMs, and defenses under different FL settings. Through the broader scope, we draw more generalizable conclusions about the attack’s effectiveness across various scenarios, demonstrating the limitations of current defenses and emphasizing the need for advanced approaches.
VII Conclusions and Future Work
@@
In this paper, we introduced PEFT-as-an-Attack (PaaA), a novel security threat to FedPEFT, demonstrating that PEFT methods can be exploited to bypass safety alignments, thus inducing language models to generate harmful content in response to malicious prompts. Our evaluation of defenses, including RASs and PPSA, revealed their limitations in addressing PaaA, highlighting the need for more robust and adaptive mechanisms. In the future, we will focus on developing advanced PPSA techniques for FedPEFT to balance the utility and safety of the language models. We also aim to integrate safety alignment directly during fine-tuning to dynamically mitigate emerging vulnerabilities without compromising model performance.
@@ Acknowledgment
The computationally intensive experiments presented in this research were performed on the Alvis [^7] https://www.c3se.chalmers.se/about/Alvis/ cluster at Chalmers University of Technology, funded by the Swedish National Infrastructure for Computing (SNIC).